# Analyzing the Impact of the Automatic Ball-Strike System in Professional Baseball: A Case Study on KBO League Data

### v1

@@ Abstract
Recent advancements in professional baseball have led to the introduction of the Automated Ball-Strike (ABS) system, or “robot umpires,” which utilize machine learning, computer vision, and precise tracking technologies to automate ball-strike calls. The Korean Baseball Organization (KBO) league became the first professional baseball league to implement ABS during the 2024 season. This study analyzes pitching data for 2,515 games across multiple KBO seasons to compare decisions made by human umpires with those made by ABS, focusing specifically on differences within the “gray zone” of the strike zone. We propose and answer four research questions to examine the differences between human and robot umpires, player adaptation to ABS, assess the ABS system’s fairness and consistency, and analyze its strategic implications for the game. Our findings offer valuable insights into the impact of technological integration in sports officiating, providing lessons relevant to future implementations in professional baseball and beyond.
@@
Sports analytics, Baseball, Automatic Ball-Strike SystemIntroduction
@@
Traditionally, umpires rely on their judgment, experience, and training to call balls and strikes. Despite rigorous training, human error leads to occasional disputes and inconsistencies [9, 10, 11]. Players sometimes exploit these inaccuracies by testing the umpire’s tendencies early in the game [12] or catchers “framing” pitches to make them appear as strikes [13, 14, 15]. ABS aims to mitigate such tactics by bringing objectivity and consistency to ball-strike calls, reducing human bias. Supporters argue that ABS ensures a fairer game, eliminating subjective errors and providing precise, consistent decisions. Additionally, ABS can enhance baseball analytics, offering insights into player performance and strategies.
@@
The adoption of the ABS system has profound implications for players, particularly pitchers and batters. For pitchers, the precise and consistent strike zone could influence pitching strategies, allowing for more predictable outcomes of their pitches. Similarly, batters can adjust their tactics at the plate, knowing that the strike zone will be consistently enforced. This change is expected to impact various aspects of the game, from pitch selection and batting averages to overall game strategy.
This research analyzes pitch data from three and a half KBO League seasons (three pre-ABS seasons and half of the ABS-adopted 2024 season) to compare human and robot umpires, focusing on the ‘gray zone’ where calls are most contentious. The analysis aims to help players and teams adapt to robot umpires. Specifically, we address four research questions:
@@
RQ2: Where does the gray zone in strike-ball calls lie and how does the strike call differ for robot umpires?
@@
The findings from this research will contribute to the discussion on integrating technology in sports officiating. Insights from the KBO’s experience with ABS will provide valuable lessons for other leagues considering similar technologies. This study enhances our understanding of how technological advancements can balance tradition and innovation in sports, shaping the future of professional baseball. By examining the ABS system’s impact on the KBO League, this research offers a comprehensive analysis of its implications and sets the direction for future exploration of technology in baseball.
@@
To compare strike-ball calls made by human and robot umpires, we compiled data from Naver Sports [21] covering 2,515 games over four years (2021 to 2024), totaling 22,934 innings. This dataset includes 941 players from 10 teams and 42 home plate umpires. While each of the ten KBO teams plays 144 games per season, discrepancies in the database led to a few missing games. We analyzed data from the 2024 season with the ABS system up to the All-Star break, and compared it with human umpire data from 2021 to 2023. Notably, the strike zone was officially expanded in 2022, resulting in a narrower zone in 2021 compared to later years. We aim to evaluate the impact of ABS adoption by comparing data from 2022-2023 and 2024. Additionally, we included 2024 season Major League Baseball pitch data, sourced from Statcast Savant [22], for further analysis.
@@
Table 1 provides an overview of our data, including detailed information on each pitch such as pitcher and batter details, pitch speed, type, location relative to the strike zone, and pitch outcome (e.g., ball, strike, foul, hit). Throughout this paper, the terms ‘robot umpire’ and ‘ABS’ are used interchangeably. The dataset will be publicly accessible at https://github.com/eis-lab/where-do-the-robot-umpires-see.
@@
In Eq 1, denotes the probability of classifying a pitch as a “strike” rather than a “ball”. is the logistic function (), the inverse of the logit function (), which maps input to a probability. affects the steepness of the probability contour, determining how the probability transitions between a strike and a ball based on the distance () from the zone’s center. Parameter defines the boundary size of the zone, influencing its width and height. The distance measure is defined as:
where and correspond to the horizontal and vertical location of a pitch measured in feet. The origin of the coordinate system is based on the ground level for the vertical axis and the home plate center for the horizontal axis. and denote the center of the zone, serving as the reference point for measuring distances. is a scaling factor that adjusts the ratio of the vertical and horizontal dimensions of the zone, while fine-tunes the curvature of the strike zone.
@@
Moreover, provides insight into the consistency of the zone: a larger indicates that pitches far from the center are less likely to be strikes, and pitches at similar distances from the center receive consistent calls. Lastly, indicates the rectilinearity of the strike zone. Analogous to the p-norm, a larger suggests a shape closer to a quadrilateral. Since the KBO’s strike zone is rectilinear, a well-modeled zone should exhibit a high .
@@ Discussion
Per-player level analysis. In this work, we primarily focused on analyzing the pitch data from a macroscopic and diachronic perspective. Consequently, in Analysis III, we did not approach our analysis from a per-player perspective. Players can react differently to the ABS system, with factors such as years of experience, batting and pitching style, historical performance metrics, and physical characteristics potentially influencing strategic changes. This would require detailed persona modeling of each player to capture these individual characteristics. We acknowledge the importance of this deeper, granular analysis and suggest it as a promising direction for future research. Understanding individualized responses to ABS will enrich our comprehension of the system’s impact and aid in refining training and adaptation strategies for players.
@@
Improving strike zone modeling. This work employs parametric logistic function-based strike zone modeling [23]. However, this method only considers pitch location projected onto a 2D plane from the home plate umpire’s perspective, limiting the detailed analysis of the ball’s actual movement. Additionally, as noted in Analysis IV, the rectilinearity parameter shows wide credible intervals due to the characteristics of the -norm function [24, 25]. We suggest that developing a more sophisticated strike zone modeling approach to address these limitations could provide a more comprehensive analysis of baseball pitch data. Such advancements would deepen our understanding of pitch dynamics and the impact of ABS, offering valuable insights for both researchers and practitioners.
@@ Conclusion
In this study, we conducted a pioneering analysis of the differences between Automated Ball-Strike (ABS) systems and human umpires, offering crucial insights for players and coaching staff. Key findings include:
Gray Zones Identified: We pinpointed “gray zones” at the strike zone’s edges where discrepancies between human and robot umpires are most pronounced, illustrating ABS’s improvements in consistency.
Pitcher and Batter Adjustments: Pitchers modified their strategies in response to ABS strike zones, while batters’ hit attempt ratios remained stable.
ABS Fairness and Consistency: ABS demonstrated greater consistency and fairness compared to human umpires, validating its potential to reduce human error and bias.
These findings provide valuable information for teams to better prepare for and adapt to ABS. Understanding the differences in strike calls and the resulting changes in game dynamics will support teams in developing effective strategies under the ABS system. Future research should include long-term, individual-level analyses and expand to other leagues like MLB and NPB to further refine strategies for integrating ABS into baseball.

### v2

# Analyzing the Impact of the Automatic Ball Strike System in Professional Baseball through a Case Study on KBO League Data

@@ Abstract
Recent advancements in professional baseball have led to the introduction of the Automated Ball-Strike (ABS) system, or “robot umpires,” which utilize machine learning, computer vision, and precise tracking technologies to automate ball-strike calls. The Korean Baseball Organization (KBO) league became the first professional baseball league to implement ABS during the 2024 season. Leveraging pitch data from 2,515 KBO games across multiple seasons and employing mathematical modeling, we examine the aggregate decision tendencies of human umpires versus those of the ABS within the “gray zone” of the strike zone. We propose and answer four research questions to examine the differences between human and robot umpires, player adaptation to ABS, assess the ABS system’s fairness and consistency, and analyze its strategic implications for the game. Our findings offer valuable insights into the impact of technological integration in sports officiating, providing lessons relevant to future implementations in professional baseball and beyond.
Keywords: Automatic ball-strike system, Baseball, Sports analytics
@@
Automatic Ball-Strike system, Baseball, Sports analyticsIntroduction
@@
Traditionally, umpires rely on their judgment, experience, and training to call balls and strikes. Despite rigorous training, human error leads to occasional disputes and inconsistencies [9, 10, 11]. Williams [11] reported that Major League Baseball umpires made 34,294 incorrect ball–strike calls (14 per game) during the 2018 season. He further attributed these errors to various human biases, including game context, umpire age, and level of experience. Players sometimes exploit these inaccuracies by testing the umpire’s tendencies early in the game [12] or catchers “framing” pitches to make them appear as strikes [13, 14, 15]. ABS aims to mitigate such tactics by bringing objectivity and consistency to ball-strike calls, reducing human bias. Supporters argue that ABS ensures a fairer game, eliminating subjective errors and providing precise, consistent decisions. Additionally, ABS can enhance baseball analytics, offering insights into player performance and strategies.
@@
However, the shift to ABS also introduces important adaptation challenges. Specifically, experienced players in the KBO reported frustration with the tighter, more uniform strike zone, necessitating changes in pitching and batting strategies [21]. For pitchers, the precise and consistent zone may lead to more predictable outcomes and influence pitch selection, while batters can adjust their tactics at the plate knowing that the strike zone will be enforced without human variability. Fans also faced a learning curve in tracking the new boundaries before coming to appreciate the system’s objectivity [22]. Nevertheless, before understanding these impacts on player performance and fan experience, it is essential to first analyze the fundamental differences between ABS and human umpiring. Accordingly, our study focuses on quantifying and characterizing divergences between ABS and human strike zones, thereby laying the groundwork for future investigations into how these technological changes affect gameplay and spectator engagement.
This research analyzes pitch data from three and a half KBO League seasons (three pre-ABS seasons and the first half of the ABS-adopted 2024 season) to compare decision patterns of human and robot umpires, with particular emphasis on the so-called “gray zone,” denoting the area in which the human umpire’s estimated strike zone and the ABS strike zone differ most significantly. By quantifying these boundary mismatches, our analysis aims to pinpoint where adaptation will be potentially most critical for players and teams. Specifically, we address four research questions:
@@
RQ2: Where does the gray zone in strike-ball calls lie, and how does the strike call differ for robot umpires?
@@
To the best of our knowledge, previous strike-zone analyses have focused primarily on biases in human umpiring or player behavior [23, 24], whereas our study uniquely examines the robot umpire’s call patterns. The findings from this research will contribute to the discussion on integrating technology in sports officiating. Insights from the KBO’s experience with ABS will provide valuable lessons for other leagues considering similar technologies. Specifically, this study aims to refine our understanding of how emerging officiating technologies might balance tradition and innovation in professional baseball. Through an examination of the ABS system’s strike-zone characteristics, we offer a preliminary assessment of potential implications and establish a foundation for future research into technology’s role in the game.
@@
To compare strike-ball calls made by human and robot umpires, we compiled data from Naver Sports [25] covering 2,515 games over four years (2021 to 2024), totaling 22,934 innings. This dataset includes 941 players from 10 teams and 42 home plate umpires. While each of the ten KBO teams plays 144 games per season, discrepancies in the database led to a few missing games. We analyzed data from the 2024 season with the ABS system up to the All-Star break, and compared it with human umpire data from 2021 to 2023. Notably, the strike zone was officially expanded in 2022, resulting in a narrower zone in 2021 compared to later years. While this change precludes direct one‐to‐one comparisons of absolute boundaries, we nonetheless include the 2021 human-umpire data to capture and characterize inherent bias and decision-making patterns that persist across seasons. We then evaluate the impact of ABS adoption by comparing data from 2022–2023 and 2024. Additionally, we included the 2024 season Major League Baseball (MLB) pitch data, sourced from Statcast Savant [26], solely for evaluating cross-league generalization as described in the Discussion section. Note that unless explicitly mentioned as the MLB data, all references to 2024 data in this paper refer to robot umpire calls in the KBO League.
@@
Table 1 provides an overview of our dataset, including the number of games and total pitches in each season, as well as counts of unique pitchers, batters, and umpiring systems (human vs. ABS). While Table 1 summarizes these high-level statistics, each individual pitch record in our data also contains detailed attributes, pitch speed, pitch type, location relative to the strike zone, pitch outcome (e.g., ball, strike, foul, hit), and context meta data (player, umpire, inning, ball count, etc), which are used for our analyses. Throughout this paper, the terms ‘robot umpire’ and ‘ABS’ are used interchangeably. The dataset will be publicly accessible at https://github.com/eis-lab/where-do-the-robot-umpires-see.
@@
In Eq 1, denotes the probability of classifying a pitch as a “strike” rather than a “ball”. is the logistic function (), which maps the input to a probability. affects the steepness of the probability contour, determining how the probability transitions between a strike and a ball based on the distance () from the zone’s center. Parameter defines the boundary size of the zone, influencing its width and height. The distance measure is defined as:
where and correspond to the horizontal and vertical location of a pitch measured in feet. The origin of the coordinate system is based on the ground level for the vertical axis and the home plate center for the horizontal axis. and denote the center of the zone, serving as the reference point for measuring distance . is a scaling factor that adjusts the ratio of the vertical and horizontal dimensions of the zone, while fine-tunes the curvature of the strike zone.
@@
Moreover, provides insight into the consistency of the zone: a larger indicates that pitches far from the center are less likely to be strikes, and pitches at similar distances from the center receive consistent calls. Lastly, indicates the rectilinearity of the strike zone. We adopt a p-norm framework to flexibly capture the zone’s geometry. The p-norm generalizes Euclidean distance (p=2) toward the Chebyshev (p) norm, allowing shapes to interpolate continuously between circular and perfectly rectangular contours. Analogous to the p-norm, a larger suggests a shape closer to a quadrilateral. Since the KBO’s strike zone is rectilinear, a well-modeled zone should exhibit a high .
We note that our strike‐zone modeling, based solely on two‐dimensional pitch coordinates and observed call labels, is a generalizable approach both for human and robot umpires. We apply this method separately to the human‐umpire data and to the ABS data for each season in our dataset to produce comparable zone estimates.
@@
Furthermore, we investigated whether human umpires and ABS exhibit systematic differences in their treatment of each pitch type by fitting a separate strike-zone model to the calls for each pitch arsenal. Figure 14 visualizes the decision boundaries of these per-arsenal models. As the plots show, while ABS deviations across types remain minimal, the human-umpire boundaries display noticeable variability from one pitch type to another. Figure 15 compares the fitted model parameters across arsenals, demonstrating quantitatively that ABS maintains a tightly clustered zone height (), width () and center point (, with significantly high consistency () and rectilinearity (), regardless of pitch type. These findings provide strong evidence that the ABS system delivers more robust and fair verdicts across different pitch arsenals than human umpires.
@@ Discussion
Potential drawback of ABS. One limitation of this study is that, by focusing exclusively on differences between human and ABS strike zones, we do not address potential drawbacks of ABS adoption, most notably, the loss of traditional umpiring dynamics and its impact on fan engagement. Removing this human element can diminish the suspense and emotional connection that close calls and umpire reactions provide, potentially detracting from the spectator experience. An interesting direction of future research is to explore how ABS influences fan perceptions and consider strategies (e.g., enhanced broadcast features or mixed‐officiating models) to preserve the game’s dynamic atmosphere.
Per-player level analysis. In this work, we adopted a macroscopic and diachronic view of ABS impact on umpiring decisions. Consequently, in Analysis III where we observed player’s adjustment to ABS, we did not perform analysis from a per-player perspective with fine-grained granularity. Note that players can react differently to the ABS system, impacted by factors such as years of experience, batting and pitching style, historical performance metrics, and physical characteristics potentially influencing strategic changes. This would require a detailed persona modeling of each player to capture these individual characteristics. We acknowledge the importance of this deeper, granular analysis and suggest it as a promising direction for future research. Over multiple seasons, for instance, pitchers may shift their pitch-selection toward locations that exploit the ABS consistency, and batters may refine their plate approach in response to a more predictable strike zone. Similarly, coaching staffs and scouts are likely to evolve training regimens and talent evaluation criteria around objective, data‐driven strike‐zone metrics. Understanding these individualized and systemic responses to ABS will enrich our comprehension of the system’s impact and aid in refining training and adaptation strategies for players.
@@
Improving strike zone modeling. This work employs parametric logistic function-based strike zone modeling [23]. However, this method only considers pitch location projected onto a 2D plane from the home plate umpire’s perspective, limiting the detailed analysis of the ball’s actual movement. Additionally, as noted in Analysis IV, the rectilinearity parameter shows wide credible intervals due to the characteristics of the -norm function [27, 28]. Moreover, our current approach applies a uniform, league-average zone height rather than adjusting for each batter’s stature, potentially biasing comparisons for very tall or short players. We suggest that developing a more sophisticated strike zone modeling approach to address these limitations could provide a more comprehensive analysis of baseball pitch data. Such advancements would deepen our understanding of pitch dynamics and the impact of ABS, offering valuable insights for both researchers and practitioners.
@@ Conclusion
In this study, we conducted a pioneering analysis of the differences between Automated Ball-Strike (ABS) systems and human umpires that move beyond prior work, which have been centered on human-umpire biases and player behavior, offering crucial insights for players and coaching staff. Key findings include:
Gray Zones Identified: We pinpointed “gray zones” at the strike zone’s edges where discrepancies between human and robot umpires are most pronounced, illustrating ABS’s improvements in consistency. We anticipate that the discrepancies in strike zones identified between human and robot umpires will meaningfully affect players’ experiences and in-game outcomes. To cope with these differences, players will need to adapt and analyze them, while league organizations should both facilitate players’ acclimation to computer-generated judgments and consider revising rules to better align with players’ established standards.
Pitcher and Batter Adjustments: Pitchers modified their strategies in response to ABS strike zones, while batters’ hit attempt ratios remained stable. These findings suggest that pitchers, who with more active options, are able to capitalize on the consistency of the ABS system, whereas batters do not yet appear to take full advantage of these predictable strike zones. Overall, this implies that batters will benefit from more specialized training, targeted adaptation, and strategic adjustments to effectively respond to ABS officiating.
ABS Fairness and Consistency: ABS demonstrated greater consistency and fairness compared to human umpires, validating its potential to reduce human error and bias. We presume that this consistent and fair strike-zone can mitigate the frustration of both players and fans due to erroneous calls.
These findings provide valuable information for teams to better prepare for and adapt to ABS. Understanding the differences in strike calls and the resulting changes in game dynamics will support teams in developing effective strategies under the ABS system. Future research should include long-term, individual-level analyses and expand to other leagues like MLB and NPB to further refine strategies for integrating ABS into baseball. In addition, more sophisticated analyses of ABS, including biases arising from various environmental conditions and external contextual factors, as well as deeper investigations using three-dimensional trajectory data, would constitute interesting and valuable future research directions.
Data availability
Data is provided within the manuscript and can be downloaded through the following URL: https://github.com/eis-lab/where-do-the-robot-umpires-see.